# weaker gravitational waves

### v2

@@ 1 Introduction
in contrast to aforementioned $t_{ann}\propto 1/V_{bias}$. Note that values of the potential bias $V_{bias}$ are typically assumed to be very (almost exponentially) small, and therefore this deviation from naive expectations can have a profound impact on GW phenomenology. Namely, for the same $V_{bias}$ one obtains a much earlier decay of the DW network compared to naive estimates, and hence a weaker GW signal, because the energy density of DWs has less time to grow relative to the dominant, e.g., radiation, energy density.
@@ 2 Theoretical setup
[…]
@@
[…]
@@ 3 Numerical setup
[…]
@@ 4 Numerical results: evolution of domain wall network
where $S$ is the DW comoving area captured within the comoving volume $V$. The numerical computation of the DW area $S$ is performed using the PRS algorithm developed in Ref. [23]. Note that in the unbiased case the parameter $\xi$ serves to diagnose when the scaling regime is established. Namely, once the scaling is reached, the parameter $\xi$ takes on a constant value. In the case $\epsilon\neq 0$, DW evolution remains approximately unaffected by the potential bias until the time $\tau\sim 5$, which reiterates the above observation made with histograms. Around the time $\tau\sim 5$, different $\xi(\tau)$-curves start to deviate from the case $\epsilon=0$. Overall, the $\xi(\tau)$-curves take bell-like shapes (for vacuum initial conditions), which are more pronounced for larger $\epsilon$. The bell width is growing upon the decrease of the bias parameter.
@@
To derive the relation (26) we have probed the ranges $\epsilon=[0.0035,0.05]$ and $\epsilon=[0.01,0.05]$ using $2048^{3}$ and $1024^{3}$ lattices, respectively. Note that the similar conclusion can be derived from the analysis of false vacuum fraction performed in Ref. [21], where the narrow range $\epsilon=[4\cdot 10^{-4},8\cdot 10^{-4}]$ has been explored with higher resolution simulations. One can check that the behavior (26) is indeed obeyed except for the smallest $\epsilon$ there. Steepening of the behavior $\tau_{ann}(\epsilon)$ in the range of very small $\epsilon$ is likely attributed to the fact that in this case the network decay takes place at sufficiently late times, when effects, related to finite lattice size/resolution, become significant.
@@
Figure 6: Comparison of the potential bias impact on DW evolution caused by the symmetry breaking potentials $V_{breaking}=\epsilon\chi^{3}$, $\epsilon\chi v^{2}$, and $\epsilon(\chi^{5}/v^{2}+\chi^{6}/v^{3})$. The bias parameter $\epsilon$ has been set to $\epsilon=0.05$ (in units of $\lambda v$). The impact is demonstrated in terms of the scaling parameter $\xi$ (left panel) and the false vacuum fraction ${\cal F}_{fv}$ (right panel). Simulations have been carried out with $1024^{3}$ lattice.
@@
Let us consider other choices of $Z_{2}$-symmetry breaking caused by the linear term $V_{breaking}=\epsilon\chi v^{2}$ as well as the non-renormalisable term of the form $V_{breaking}=\epsilon\chi^{5}/v^{2}+\epsilon^{\prime}\chi^{6}/v^{3}$. In the latter case, the term $\epsilon^{\prime}\chi^{6}/v^{3}$ with some positive constant $\epsilon^{\prime}$ has been added to keep the overall potential bounded from below. As it is clear from Fig. 6, the impact of the higher order potential bias on DW evolution is almost indistinguishable from that of the cubic bias. On the other hand, the effect of the linear bias is dramatically different. The likely reason is that the linear bias term changes behavior of its potential $V(\chi)$ around the minimum $\chi=0$ leading to non-zero $\partial V/\partial\chi$ at $\chi=0$. As a result, one of the minima becomes slightly preferred meaning that the linear potential bias triggers the population bias. This may explain, why the DW network is destroyed at much faster rate for the same $\epsilon$ [34], and consequently the expected GW signal is suppressed relative to the case of the cubic bias. In this work we focus on the effects, which are entirely due to the potential bias, and omit any further discussion of the case $V_{breaking}\propto\chi$.
@@ 6 Conclusions
From the viewpoint of observations, the most interesting piece of information is contained in the GW power spectrum shape. While the spectral shape of GWs produced by the network of biased walls is similar to that in the case $\epsilon=0$, there are important differences most notably in the UV part, where GWs from biased DWs exhibit larger power compared to the unbiased case. In this regard, we confirm the results obtained recently in Ref. [17], which uses a different code, but disagree with those in Ref. [18], which exploits a modified CosmoLattice code. This tension is worth exploring in the future. In terms of the fitting formula (33), we have restricted values of the parameters $a,~{}b$ and $c$. Note that the plateau part of the spectrum observed in Refs. [27, 28] for unbiased DWs is also present in the biased case, cf. Ref. [16]. Interestingly, in terms of conformal momentum, the position of IR edge of the plateau exhibits a remarkable stability, i.e., it remains constant as a function of time and potential bias. This may suggest that the plateau, if due to some physical mechanism rather than a simulation artefact, starts at $k/a\sim 4\pi\sqrt{H/\delta_{wall}}$. While there is currently no consensus on the nature of the plateau, our simulations in this paper and in Ref. [28] indeed suggest that the plateau has a physical origin.

### v3

@@ 1 Introduction
in contrast to aforementioned $t_{ann}\propto 1/V_{bias}$. Note that values of the potential bias $V_{bias}$ are typically assumed to be very (almost exponentially) small, and therefore this deviation from naive expectations can have a profound impact on GW phenomenology. Namely, for the same $V_{bias}$ one obtains a much earlier decay of the DW network compared to naive estimates, and hence a weaker GW signal, because the energy density of DWs has less time to grow relative to the dominant, e.g., radiation, energy density. Note that the result (1) has been derived for not extremely small $V_{bias}$, see Sec. 4. This has a consequence that the time span of simulations prior to the DW annihilation is rather limited, and there is no enough time for the scaling regime to be properly established. In this situation, the evolution can be sensitive to the choice of initial conditions. We address these shortcomings in Sec. 4 by varying initial conditions and by including the data obtained in Ref. [21], where higher resolution simulations allowed to probe much smaller values of $V_{bias}$ compared to our case.
@@ 2 Theoretical setup
where $a(t)$ is the Universe scale factor. Throughout this work, we assume that DWs live in the radiation-dominated Universe with the scale factor growing as $a(t)\propto\sqrt{t}$. Below we often make use of the conformal time $\tau$ defined as $\tau\sim\int dt/a(t)$, so that $a(\tau)\propto\tau$.
@@
The latter is defined as the integral over the DW profile (kink), i.e., $\sigma_{wall}=\int^{+\infty}_{-\infty}dz^{\prime}T_{00}(z^{\prime})$, where $T_{00}(z)$ is the 0000the component of the field $\chi$ stress-energy tensor. As it follows from Eq. (7), the energy density of DWs grows fast relative to that of radiation, i.e., $\rho_{wall}/\rho_{rad}\propto a^{2}$, where $a$ is the Universe scale factor. This leads to the DW problem and motivates the introduction of the potential bias.
@@ 3 Numerical setup
$\kappa^{\prime}=\pi\kappa~\mbox{(IR)}\,,\qquad\kappa^{\prime}=\frac{\pi\kappa}{2}~\mbox{(middle)}\,,\qquad\kappa^{\prime}=\frac{\pi\kappa}{6}~\mbox{(UV)}\;.$ (23)
@@ 4 Numerical results: evolution of domain wall network
where $S$ is the DW comoving area captured within the comoving volume $V$. The numerical computation of the DW area $S$ is performed using the estimator developed in Ref. [23]. Note that in the unbiased case the parameter $\xi$ serves to diagnose when the scaling regime is established. Namely, once the scaling is reached, the parameter $\xi$ takes on a constant value. In the case $\epsilon\neq 0$, DW evolution remains approximately unaffected by the potential bias until the time $\tau\sim 5$, which reiterates the above observation made with histograms. Around the time $\tau\sim 5$, different $\xi(\tau)$-curves start to deviate from the case $\epsilon=0$. Overall, the $\xi(\tau)$-curves take bell-like shapes (for vacuum initial conditions), which are more pronounced for larger $\epsilon$. The bell width is growing upon the decrease of the bias parameter.
@@
$\epsilon$ Lattice Reals $k_{cut}$ $\chi^{2}$, $\alpha=1/2$ $\chi^{2}$, $\alpha=1/3$ $\alpha$, best fit
$[0.01,0.05]$ $024^{3} 11$ $0.3$ $2.2$ $0.05$ $0.3$
$[0.01,0.05]$ $1024^{3}$ 44 11 $23.1$ $0.29$ $0.34$
$[0.01,0.05]$ $024^{3} 11$ 55 $4.3$ $0.7$ $0.34$
$[0.0035,0.05]$ $1024^{3}$ 44 11 $8.6$ $1.1$ $0.38$
$[0.0035,0.05]$ $2048^{3}$ 11 11 $42.2$ $1.9$ $0.37$
$[0.0009,0.0016]$ $3240^{3}$ 44 - $5.0$ $0.05$ $0.32$
$[0.0008,0.0016]$ $3240^{3}$ 44 - $3.0$ $1.1$ $0.42$
Table 1: The power-law behaviour $\tau_{ann}\propto 1/\epsilon^{\alpha}$ has been studied using the $\chi^{2}$ analysis for the cases $\alpha=1/2$ and $\alpha=1/3$. The results are demonstrated for a selection of lattices at different resolutions/for different initial conditions of the field $\chi$/for different number of realizations of the initial scalar field configuration. The best fit value $\alpha$ is also shown in the Table. Numbers in the last two lines have been filled using the results of simulation data from Ref. [21]. The cutoff scale on vacuum initial conditions is not specified there.
@@
To derive the relation (26) we have probed the ranges $\epsilon=[0.0035,0.05]$ and $\epsilon=[0.01,0.05]$ using $2048^{3}$ and $1024^{3}$ lattices, respectively. We have performed $\chi^{2}$-analysis to test the hypothesis $\tau_{ann}\propto 1/\epsilon^{1/3}$ versus $\tau_{ann}\propto 1/\epsilon^{1/2}$. Results shown in Table 1 demonstrate a clear preference for the former hypothesis over the latter one. Steepening of the behavior $\tau_{ann}(\epsilon)$ in the range of very small $\epsilon$ is likely attributed to the fact that in this case the network decay takes place at sufficiently late times, when effects, related to finite lattice size/resolution, become significant. On the other hand, in this range the outcome is less dependent on the possible impact of initial conditions, and therefore the steepening may indicate instead that the commonly assumed law $\tau_{ann}\propto 1/\epsilon^{1/2}$ is restored at very small $\epsilon$.
To understand this issue better, we have analysed the results of numerical simulations for the false vacuum decay from Ref. [21], where the range $\epsilon=[0.0008,0.0016]$ has been explored with higher resolution simulations. The corresponding $\chi^{2}$-values are also demonstrated in Table 1. We again observe preference of the behaviour $\tau_{ann}\propto 1/\epsilon^{1/3}$ over $\tau_{ann}\propto 1/\epsilon^{1/2}$. Simulations of Ref. [21] also exhibit a significant steepening of the behaviour $\tau_{ann}(\epsilon)$, if one allows for large simulation times, as in our case. As the ranges of $\epsilon$ considered in this work and in Ref. [21] do not overlap, this feature is perhaps a numerical artefact. Nevertheless, it serves as a motivation for further studies of the behaviour $\tau_{ann}(\epsilon)$ using high resolution simulations and for a larger scope of initial conditions.
@@
Figure 6: Comparison of the potential bias impact on DW formation and evolution caused by the symmetry breaking potentials $V_{breaking}=\epsilon\chi^{3}$, $\epsilon\chi v^{2}$, and $\epsilon(\chi^{5}/v^{2}+\chi^{6}/v^{3})$. The bias parameter $\epsilon$ has been set to $\epsilon=0.05$ (in units of $\lambda v$). The impact is demonstrated in terms of the scaling parameter $\xi$ (left panel) and the false vacuum fraction ${\cal F}_{fv}$ (right panel). Simulations have been carried out with $1024^{3}$ lattice. DWs are not formed in the case of a linear potential bias (for the choice of $\epsilon$ assumed).
@@
Let us consider other choices of $Z_{2}$-symmetry breaking caused by the linear term $V_{breaking}=\epsilon\chi v^{2}$ as well as the non-renormalisable term of the form $V_{breaking}=\epsilon\chi^{5}/v^{2}+\epsilon^{\prime}\chi^{6}/v^{3}$. In the latter case, the term $\epsilon^{\prime}\chi^{6}/v^{3}$ with some positive constant $\epsilon^{\prime}$ has been added to keep the overall potential bounded from below. As it is clear from Fig. 6, the impact of the higher order potential bias on DW evolution is almost indistinguishable from that of the cubic bias. On the other hand, the effect of the linear bias is dramatically different. The likely reason is that the linear bias term changes behavior of its potential $V(\chi)$ around the minimum $\chi=0$ leading to non-zero $\partial V/\partial\chi$ at $\chi=0$. As a result, one of the minima becomes slightly preferred meaning that the linear potential bias triggers the population bias. This may explain, why the DW network is not formed in this case, cf. Ref. [34]. In this work we focus on the effects, which are entirely due to the potential bias, and omit any further discussion of the case $V_{breaking}\propto\chi$.
@@ 6 Conclusions
It is worth stressing that our results regarding the DW annihilation time, while being at odds with naive theoretical expectations, do not contradict the numerical data obtained in the literature. Indeed, in Sec. 4 we have found that our results are in a comfortable agreement with the results of numerical simulations of Ref. [21]. Notably, the latter deals with particularly small $\epsilon$ values, for which the system reaches scaling with a high accuracy, and hence one can ignore the possible impact of initial conditions. However, in Ref. [21] the standard behaviour $t_{ann}\propto 1/\epsilon$ has been assumed. This is justified, if one sticks to a sufficiently narrow range of $\epsilon$, but discrepancy from the law $t_{ann}\propto 1/\epsilon^{2/3}$ becomes dramatic, once the extrapolation to tiny $\epsilon$ is considered.
From the viewpoint of observations, the most interesting piece of information is contained in the GW power spectrum shape. While the spectral shape of GWs produced by the network of biased walls is similar to that in the case $\epsilon=0$, there are important differences most notably in the UV part, where GWs from biased DWs exhibit larger power compared to the unbiased case. In this regard, we confirm the results obtained recently in Ref. [17], which uses a different code, but disagree with those in Ref. [18], which exploits a modified CosmoLattice code. This tension is worth exploring in the future. In terms of the fitting formula (33), we have restricted values of the parameters a,ba,~b and $c$. Note that the plateau part of the spectrum observed in Refs. [27, 28] for unbiased DWs is also present in the biased case, cf. Ref. [16]. Interestingly, in terms of conformal momentum, the position of IR edge of the plateau exhibits a remarkable stability, i.e., it remains constant as a function of time and potential bias. This may suggest that the plateau, if due to some physical mechanism rather than a simulation artefact, starts at $k/a\sim 4\pi\sqrt{H/\delta_{wall}}$. While there is currently no consensus on the nature of the plateau, our simulations in this paper and in Ref. [28] indeed suggest that the plateau has a physical origin.